(D^2 - D'^2 - 3D + 3D')z = xy


1
Expert's answer
2021-02-01T19:07:07-0500

Given equation is (D^2 - D'^2 - 3D + 3D')z = xy

Auxilliary equation is "m^2-1-3m+3 = 0"

"m=1,2"


Then, "z = f_1(x+y)+f_2(3x+y)"


C.F. "\frac{1}{(D^2 - D'^2 - 3D + 3D')}xy = \frac{1}{3D'(1+\frac{D^2 - D'^2 - 3D}{3D'})}xy"

"\frac{1}{3D'}(1+\frac{D^2 - D'^2 - 3D}{3D'})^{-1}xy =\frac{1}{3D'}(1-\frac{D^2 - D'^2 - 3D}{3D'})xy"

"= \frac{1}{3D'}[xy-\frac{1}{3D'}(-3y)] =\frac{1}{3D'}( xy+\frac{y^2}{2}) = \frac{1}{6}xy^2 + \frac{1}{18}y^3"


So, the solution of the equation is "z = f_1(x+y)+f_2(3x+y)+\frac{1}{6}xy^2 + \frac{1}{18}y^3"